John has 7 more pennies than nickels and 2 fewer quarters than pennies. His total amount of money is $2.87. How many of each type of coin does he have?
Answer by stanbon!(97)

Let number of nickels be "x".
Then number of pennies is "x+7"
And number of quarters is "x+7-2" = "x+5"
Value of nickels is (5)x cents
Value of pennies is (1)(x+7)= x+7 cents
Value of quarters is (25)(x+5)= 25x+125 cents
EQUATION:
Value of nickels + Value of pennies + Value of quarters = 287 cents
5x + x+7 + 25x+125 = 287
31x = 155
x = 5 (number of nickels)
x+7 = 12 (number of pennies)
x+5 = 10 (number of quarters)